OpenStudy (anonymous):

integral of sin^3(3x)*cos^5(3x)dx

OpenStudy (zarkon):

\[\sin^3(3x)\cos^5(3x)=\sin(3x)\sin^2(3x)\cos^5(3x)=\sin(3x)(1-\cos^2(3x))\cos^5(3x)\]

OpenStudy (anonymous):

wow youre amazing!

OpenStudy (anonymous):

but there is more to it right?

OpenStudy (zarkon):

now use u-substitution

OpenStudy (anonymous):

so u =sin(3x)?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

then it equals the cos?

OpenStudy (zarkon):

yes u=cos(3x)

OpenStudy (anonymous):

so du= -sin(3x) ! :)

OpenStudy (zarkon):

no...but close

OpenStudy (zarkon):

chain rule

OpenStudy (anonymous):

times 3 also!

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

and don't forget your dx

OpenStudy (anonymous):

and now dv=sin(3x)?

OpenStudy (anonymous):

right! youre so amazing!

OpenStudy (zarkon):

please stop saying that.... \[du= -\sin(3x)3 dx\]

OpenStudy (anonymous):

yes thats what I got

OpenStudy (anonymous):

im sorry

OpenStudy (zarkon):

now replace everything, simplify (or expand) and then integrate

OpenStudy (anonymous):

dont I need to also get dv?

OpenStudy (zarkon):

dv? what is v are you thinking about integration by parts?

OpenStudy (anonymous):

oh yeah that must be it... so in u subsitution I just use u and then replace it?

OpenStudy (zarkon):

just u

OpenStudy (zarkon):

find du...replace the stuff...then integrate

OpenStudy (anonymous):

how can i replace the sin(3x) though and the 1-cos^2

OpenStudy (anonymous):

cuz I dont have those in either u nor du

OpenStudy (zarkon):

\[du=-\sin(3x)3dx\] \[-\frac{1}{3}du=\sin(3x)dx\]

OpenStudy (anonymous):

I have to go, but thank you so much for your help. I really appreciate it. I have to look into this u subsitution some more.

OpenStudy (anonymous):

oh I see.. so change the du accordingly...thank you a lot

OpenStudy (zarkon):

\[1-\cos^2(3x)\] becomes \[1-u^2\]